OpenStudy (anonymous):

find the integral of (2x2 -x)/ (sqrt)x

zepdrix (zepdrix):

Step one: \[\Large\bf\sf \int\limits \frac{2x^2-x}{x^{1/2}}\;dx\quad=\quad \int\limits \frac{2x^2}{x^{1/2}}-\frac{x}{x^{1/2}}\;dx\]

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits 2x^{2-1/2}-x^{1-1/2}\;dx\]Remember how to divide terms of similar bases? Understand this step?

OpenStudy (anonymous):

yes... I got an answer like X^(3/2){4/5x -2/3}

zepdrix (zepdrix):

Mmm ya looks good! :)

OpenStudy (anonymous):

is it right? :)

zepdrix (zepdrix):

i think so :3 i'll double check i guess lol

OpenStudy (anonymous):

I worked from your first step and it was absolutely fine. Thanks @zepdrix

zepdrix (zepdrix):

cool, good job!

OpenStudy (anonymous):

Thank you, I'm revising now, so I might be asking more questions. Wana keep a lookout?

zepdrix (zepdrix):

Sure. I have a bunch of homework to catch up on, but I'll look up at the site from time to time. Plenty of smarty pants around if I'm not available though, like mr penguin there XD

OpenStudy (anonymous):

excellent. Thank you

OpenStudy (loser66):

cannot access to the net, hehehe

OpenStudy (loser66):

what I have is Uh-oh, lost the connection

OpenStudy (anonymous):

you're a funny guy